OpenStudy (anonymous):

:( im still lost... here's the question lim x→−2+ (x/x+1)(2x+5)/(x 2 +x) supposd to find the limit algebraically. There is no graph.

OpenStudy (anonymous):

i am a bit stumped as to what this says

OpenStudy (anonymous):

\[\frac{x}{x+1}\times\frac{2x+5}{x^2+x}\]?

OpenStudy (anonymous):

maybe if you write it again, i can figure it out. i know we want the limit as x -> -2 from the right. i just don't know what you are trying to take the limit of

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

is there an equation that i use? or do I have to grind out numbers approaching -2 from the right?